Apply the fundamental theorem of homomorphism to prove that

R^4/R^2 is isomorphic to R^2


1
Expert's answer
2021-02-25T18:29:27-0500

Consider the groups "(\mathbb R^4, +)" and "(\mathbb R^2, +)", and define the map "\psi: \mathbb R^4 \to\mathbb R^2, \ \psi(x_1,x_2,x_3,x_4)=(x_1,x_2)."


Since "\psi((x_1,x_2,x_3,x_4)+(y_1,y_2,y_3,y_4))=\psi(x_1+y_1,x_2+y_2,x_3+y_3,x_4+y_4)=
(x_1+y_1, x_2+y_2)=(x_1,y_1)+(x_2,y_2)=\psi(x_1,x_2,x_3,x_4)+\psi(y_1,y_2,y_3,y_4),"

we conclude that "\psi" is a homomorphism.


Let us find kernel:


"\ker(\psi)=\{(x_1,x_2,x_3,x_4)\in\mathbb R^4\ |\ \psi(x_1,x_2,x_3,x_4)=(0,0)\} =
\{(x_1,x_2,x_3,x_4)\in\mathbb R^4\ |\ (x_1,x_2)=(0,0)\} 
=\{(0,0,x_3,x_4)\ |\ x_3,x_4\in\mathbb R\}."


The bijective map "f: \mathbb R^2 \to \ker(\psi),\ f(a_1,a_2)=(0,0,a_1,a_2)," is an isomomorphism. Indeed, "f((a_1,a_2)+(b_1,b_2))=f(a_1+b_1,a_2+b_2)=\n(0,0,a_1+b_1, a_2+b_2)=(0,0,a_1,a_2)+(0,0,b_1,b_2)=f(a_1,a_2)+f(b_1,b_2)."


Taking into account that "Im(\psi)=\mathbb R^2,\ \ker(\psi)\cong\mathbb R^2," and according to  the fundamental theorem of homomorphism "\mathbb R^4/\ker(\psi)\cong Im(\psi)", we conclude that "\mathbb R^4/\mathbb R^2\cong \mathbb R^2."